It takes 2664 J of energy to heat 80.0 g of substance A from 10.0oC to 25.0oC. What is the specific heat of substance A in J/goC? Round your answer to three sig figs.
1
Expert's answer
2020-05-09T14:12:07-0400

"q=cm(T_2-T_1)"

"c= \frac{q}{m(T_2)-T_1}= \frac{2664 J}{80.0 g(25.0-10.0 ^\circ C)}= 2.22 \frac{J}{g\times ^\circ C}"